# FlowDAS: A Flow-Based Framework for Data Assimilation

### v1

@@ Abstract
Data assimilation (DA) is crucial for improving the accuracy of state estimation in complex dynamical systems by integrating observational data with physical models. Traditional solutions rely on either pure model-driven approaches, such as Bayesian filters that struggle with nonlinearity, or data-driven methods using deep learning priors, which often lack generalizability and physical interpretability. Recently, score-based DA methods have been introduced, focusing on learning prior distributions but neglecting explicit state transition dynamics, leading to limited accuracy improvements. To tackle the challenge, we introduce FlowDAS, a novel generative model-based framework using the stochastic interpolants to unify the learning of state transition dynamics and generative priors. FlowDAS achieves stable and observation-consistent inference by initializing from proximal previous states, mitigating the instability seen in score-based methods. Our extensive experiments demonstrate FlowDAS’s superior performance on various benchmarks, from the Lorenz system to high-dimensional fluid super-resolution tasks. FlowDAS also demonstrates improved tracking accuracy on practical Particle Image Velocimetry (PIV) task, showcasing its effectiveness in complex flow field reconstruction.
@@ 1 Introduction
Recovering state variables in complex dynamical systems is a fundamental problem in many scientific and engineering domains. Accurate state estimation from noisy and incomplete data is essential in applications like climate modeling, weather forecasting, and seismology to understand underlying physical processes and make reliable predictions. For example, in fluid dynamics, we want to recover a continuous velocity field from sparse, noisy observations, where the task is governed by nonlinear, time-dependent partial differential equations (PDEs). The problem is challenging to solve due to the inherent stochasticity and high dimensionality of the processes. Mathematically, a discrete-time stochastic dynamical system can be characterized by the state-space model:
@@
where $\bm{x}_{k}\in\mathbb{R}^{D}$ is the state vector at time step $k$, $\Psi(\cdot)$ is the state transition map, and $\bm{\xi}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I}_{D})$ denotes Gaussian noise. The observations $\bm{y}_{k}\in\mathbb{R}^{M}$ are related to the state through the measurement map $\mathcal{A}(\cdot)$, with observation noise $\bm{\eta}_{k}\sim\mathcal{N}(\mathbf{0},\gamma^{2}\bm{I}_{M})$. Our goal is to estimate the posterior of the state trajectory $\bm{x}_{1:K}$ given observations $\bm{y}_{1:K}$ and the initial state $\bm{x}_{0}$, i.e., $p(\bm{x}_{1:K}\mid\bm{y}_{1:K},\bm{x}_{0})$, as shown in Figure 1.
This estimation problem, commonly referred to as Data Assimilation (DA) (Lahoz and Menard, 2010; Law et al., 2015; Bouttier and Courtier, 2002), originated in atmospheric (Wang et al., 2000; Bocquet et al., 2015; Reichle, 2008) and oceanic forecasting (Derber and Rosati, 1989; Cummings, 2005; Cummings and Smedstad, 2013; Carton et al., 2000; Bertino et al., 2003), and has recently found numerous applications across diverse scientific and engineering domains (Rabier, 2005; Geer et al., 2018; Gustafsson et al., 2018; Rodell et al., 2004; Van Leeuwen, 2010; Fletcher, 2017; Carrassi et al., 2018). DA combines observations with model predictions to provide more accurate and physically consistent state estimates. Despite significant advancements, existing DA methods still face significant challenges due to high nonlinearity and high dimensionality of the dynamic processes.
Existing DA techniques can be divided into model-driven and data-driven approaches. Model-driven methods, like Bayesian filters, rely on explicit physical models to describe dynamics for estimating the states. For example, Kalman filters and their variants (e.g., extended, unscented, and ensemble Kalman filters) (Evensen, 2003) assume (quasi-)linear state transitions and observations, while particle filters (Gordon et al., 1993) offer more flexibility by using Monte Carlo sampling for Bayesian estimation. However, model-driven methods face the curse of dimensionality in real-world applications (Bickel et al., 2008), because they typically require costly numerical solvers for explicit physical modeling. In contrast, data-driven approaches leverage machine learning methods to directly learn state dynamics from observational data, bypassing the usage of explicit physical models. For example, recent developments in score-based diffusion models demonstrate promise for capturing complex state transitions by learning from sequences of dynamic states (Qu et al., 2024; Rozet and Louppe, 2023). These generative models approximate the joint distribution of entire state sequences, akin to popular video generation models, and can be employed as generative priors within a Bayesian framework to support DA. However, these models often encode dynamics in latent spaces that lack accuracy and physical interpretability, leading to instability during inference and yielding biased state estimates.
Figure 1: An overview of FlowDAS. We introduce a flow-based framework for data assimilation, named FlowDAS, to estimate the state trajectory $\bm{x}_{0:K}$ from the noisy observations $\bm{y}_{1:K}$. Our framework uses a flow-based stochastic differential equation to model the stochastic dynamics of the system and leverages the observations to improve the prediction accuracy. On the right, we show a comparison of Navier-Stokes flow results between FlowDAS and the Score-based Data Assimilation (Rozet and Louppe, 2023) method. FlowDAS produces more accurate prediction compared to the score-based method, demonstrating its strong capability for data assimilation of high-dimensional complex systems.
To overcome these limitations, we propose FlowDAS: a flow-based framework for data assimilation. FlowDAS utilizes stochastic interpolants (Albergo et al., 2023), a new class of flow-based generative models, to learn neural surrogates of complex system dynamics. Stochastic interpolant models employ SDEs to create interpolative pathways between arbitrary data distributions, providing flexible control over transition dynamics for smooth, stable generation that is well-suited for capturing nonlinear dynamics in DA tasks. This learned flow model is then integrated with observational data to enable accurate and efficient state estimation. Extensive experiments demonstrate that FlowDAS achieves state-of-the-art performance across diverse scientific applications, including the chaotic Lorenz system and Navier-Stokes equations in fluid mechanics. We further validate our approach in a practical DA application, Particle Image Velocimetry (PIV) (Raffel et al., 2018), where it effectively estimates dense fluid velocity maps from sparse particle motion observations.
2 Related Work
2.1 Existing Data Assimilation Approaches
DA approaches can be broadly categorized into model-driven and data-driven methods, with particle filters and score-based data assimilation (SDA) as representative approaches in each type.
Particle filters
Gordon et al. (1993) are a class of sequential Monte Carlo (SMC) methods commonly used for DA in non-linear, non-Gaussian dynamical systems. They represent the state distribution as a set of weighted particles, each acting as a hypothesis for the state of the system. As new observations are received, particle states are updated on the basis of the dynamics of the system in Equation 1, and the particle weights are adjusted according to the likelihood of each state given the observation. Low-weight particles are periodically resampled to preserve computational resources on the most probable states. Particle filters are highly flexible, able to model complex distributions without requiring linearity or Gaussian assumptions. However, they face significant computational challenges in high-dimensional settings due to the curse of dimensionality (Bickel et al., 2008), limiting their scalability in large physical systems.
Score-based data assimilation
@@ 2.1 Existing Data Assimilation Approaches
(SDA) (Rozet and Louppe, 2023) is a recent data-driven approach that employs score-based diffusion models to estimate state trajectories in dynamical systems. It bypasses explicit physical modeling by learning the joint distribution of short state trajectory segments (e.g., $12k+$ 1 time steps) via a score network $\bm{s}_{\theta}(\bm{x}_{i-k:i+k})$. By integrating the score network with the observation model in a diffusion posterior sampling (DPS) framework (Chung et al., 2022), SDA can generate entire state trajectories in zero-shot and non-autoregressive manners. This approach is computationally efficient for high-dimensional systems, as generative priors on short segments allow for parallel inference, making it particularly useful for systems with sparse or limited observations. Despite these advantages, because the latent dynamics may fail to capture real-world physics, the learned dynamics of SDA may lack physical interpretability, leading to potential inaccuracies. SDA also requires substantial training data, and its posterior approximations may be less reliable in systems where the physical model is highly sensitive, as illustrated by the double well potential problem in Section B.2.
@@ 2.2 Stochastic Interpolants
Stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023; Chen et al., 2024b) is a recent generative modeling framework that unifies flow-based and diffusion-based models. It provides a smooth, controlled transition between arbitrary probability densities over a finite time horizon.
Consider a stochastic process $\bm{X}_{s}$ defined over the interval $s\in[0,1]$, which evolves from an initial state $\bm{X}_{0}\sim\pi(\bm{X}_{0})$ to the final state $\bm{X}_{1}\sim q(\bm{X}_{1})$. This process defines a smooth path from the base distribution to the target distribution, facilitating generative modeling. A stochastic interpolant can be described as (Chen et al., 2024b):
@@
where $(\bm{X}_{0},\bm{X}_{1})\sim p(\bm{X}_{0},\bm{X}_{1})$. $\bm{W}_{s}$ is a Wiener process for $s\in[0,1]$ introduced after $\bm{X}_{0}$ and $\bm{X}_{1}$ are sampled, ensuring that $\bm{W}_{s}$ is independent of $\bm{X}_{0}$ and $\bm{X}_{1}$. The time-varying coefficients $\alpha_{s},\beta_{s},\sigma_{s}\in C^{1}([0,1])$ satisfy boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=\sigma_{1}=0$. A typical choice for these coefficients is $\alpha_{s}=1-s$, $\sigma_{s}=1-s$, and $\beta_{s}=s^{2}$, ensuring that $\tilde{\bm{I}}_{0}=\bm{X}_{0}$ and $\tilde{\bm{I}}_{1}=\bm{X}_{1}$, thereby creating a smooth interpolation from $\bm{X}_{0}$ to $\bm{X}_{1}$. Moreover, Chen et al. (2024b) further showed that, for all $(s,\bm{X}_{0})\in[0,1]\times\mathbb{R}^{D}$, $\tilde{\bm{I}}_{s}\mid\bm{X}_{0}$ has the same distribution as $\bm{X}_{s}$, which is the solution to the following SDE:
@@
The “velocity" of the interpolant path, $\bm{R}_{s}$, is given by:
$\bm{R}_{s}=\dot{\alpha}_{s}\bm{X}_{0}+\dot{\beta}_{s}\bm{X}_{1}+\dot{\sigma}_{s}\bm{W}_{s}.$ (6)
This formulation captures the instantaneous rate of change along the interpolation path, ensuring smooth transitions that are well-suited for modeling complex distributions.
Furthermore, the score function $\nabla_{\bm{X}_{s}}\log p(\bm{X}_{s}\mid\bm{X}_{0})$ for this SDE can be expressed as:
$\nabla\log p(\bm{X}_{s}\mid\bm{X}_{0})=\lambda_{s}\left[\beta_{s}\bm{b}_{s}(\bm{X}_{s},\bm{X}_{0})-\bm{c}_{s}(\bm{X}_{s},\bm{X}_{0})\right],$ (7)
where $\lambda_{s}$ and $\bm{c}_{s}(\bm{X}_{s},\bm{X}_{0})$ are defined to control the score-based dynamics.
Stochastic interpolants provides a robust method for estimating the conditional distribution $p(\bm{x}_{k+1}\mid\bm{x}_{k})$, which offers a powerful surrogate state transition model. In DA, given the current state $\bm{x}_{k}=\bm{X}_{0}$, this framework can evolve the process $\bm{X}_{s}$ smoothly from $s=0$ to $s=1$, to capture the transition dynamics to $\bm{x}_{k+1}$. However, achieving observation-consistent predictions requires conditioning these transitions on observational data. The next section introduces our approach, which leverages stochastic interpolants for observation-informed state prediction, enabling more accurate and reliable DA.
3 Methods
This section presents our approach, which extends the original stochastic interpolants framework to incorporate observational data for effective DA. Inspired by DPS (Chung et al., 2022) and related works (Li et al., 2024; Alkhouri et al., 2024; Song et al., 2024), our method adapts the score function into a conditional form, aligning data-driven state transition modeling with observed data to enable accurate, observation-consistent state estimation in complex dynamical systems.

### v2

# FlowDAS: A Stochastic Interpolant-based Framework for Data Assimilation

@@ Abstract
Data assimilation (DA) integrates observations with a dynamical model to estimate states of PDE-governed systems. Model-driven methods (e.g., Kalman, particle) presuppose full knowledge of the true dynamics, which is not always satisfied in practice, while purely data-driven solvers learn a deterministic mapping between observations and states and therefore miss the intrinsic stochasticity of real processes. Recently, Score-based diffusion models learn a global diffusion prior and provide a good modelling of the stochastic dynamics, showing new potential for DA. However, their all-at-once generation rather than the step-by-step transition limits their performance when dealing with highly complex stochastic processes and lacks physical interpretability. To tackle these drawbacks, we introduce FlowDAS, a generative DA framework that uses stochastic interpolants to directly learn state transition dynamics and achieve step-by-step transition to better model the real dynamics. We also improve the framework by combining the observation, better suiting the DA settings. Directly learning the underlying dynamics from collected data removes restrictive dynamical assumptions, and conditioning on observations at each interpolation step yields stable, measurement-consistent forecasts. Experiments on Lorenz-63, Navier–Stokes super-resolution/sparse-observation scenarios, and large-scale weather forecasting—–where dynamics are partly or wholly unknown—show that FlowDAS surpasses model-driven methods, neural operators, and score-based baselines in accuracy and physical plausibility. The source code is available at https://github.com/umjiayx/FlowDAS.git.
@@ 1 Introduction
Recovering state variables in complex dynamical systems is a fundamental problem in science and engineering. Accurate state estimation from noisy, incomplete data is critical in weather forecasting [10, 42], oceanography [18, 19], seismology [63, 5], and many other fields [61], where reliable predictions depend on understanding the underlying physics [9, 6, 41]. A representative example is fluid dynamics [53, 11, 20], where one aims to reconstruct a continuous velocity field from sparse, noisy observations governed by nonlinear, time-dependent partial differential equations (PDEs)—a task complicated by stochasticity and high dimensionality. To meet these challenges, data assimilation (DA) combines model forecasts with observations to produce physically consistent state estimates; developed first in atmospheric and oceanic forecasting, DA is now ubiquitous across many scientific and engineering domains [40, 25, 27, 44, 56, 23, 12].
Mathematically, a discrete-time stochastic dynamical system can be described by:
@@
where $\bm{x}_{k}\in\mathbb{R}^{D}$ is the state vector at time step $k$, $\Psi(\cdot)$ is the state transition map, and $\bm{\xi}_{k}\in\mathbb{R}^{D}$ denotes the stochastic force. The observations $\bm{y}_{k}\in\mathbb{R}^{M}$ are related to the state through the measurement map $\mathcal{A}(\cdot)$, with observation noise $\bm{\eta}_{k}\sim\mathcal{N}(\mathbf{0},\gamma^{2}\bm{I}_{M})$. In many practical settings we have an initial window of known states $\bm{x}_{0:L-1}$, after which observations $\bm{y}_{L:K}$ will only be noisy and partial. DA seeks to infer the posterior of the state trajectory $\bm{x}_{L:K}$ given observations $\bm{y}_{L:K}$ and the initial states $\bm{x}_{0:L-1}$, i.e., $p(\bm{x}_{L:K}\mid\bm{y}_{L:K},\bm{x}_{0:L-1})$, as shown in Figure 1. Moreover, when the observations are absent, the task reduces to probabilistic forecasting [49], where we predict $p(\bm{x}_{L:K}\mid\bm{x}_{0:L-1})$.
Figure 1: An overview of FlowDAS. We introduce a stochastic interpolant-based framework for data assimilation, named FlowDAS, to estimate states $\bm{x}_{L:K}$ from the noisy (sparse or low-resolution) observations $\bm{y}_{L:K}$. FlowDAS models the stochastic dynamics of the system with a flow-based stochastic differential equation (SDE) and incorporates the observations to improve the prediction accuracy. On the right, we show a conceptual illustration of the geometry of the process to estimate $\hat{\bm{x}}_{k+1}$ from $\bm{x}_{k}$. $\mathcal{M}_{s}$ denotes the generative manifold at interpolation step $s$. The gradient guidance $\nabla_{\bm{x}_{s}}\|\bm{y}-\mathcal{A}(\hat{\bm{x}})\|_{2}^{2}$ enforces observation consistency during the generation process.
Existing DA methods split into two categories. Model-driven methods—Kalman variants for quasi-linear/Gaussian systems [22] and particle filters for fully nonlinear cases [26]—need an accurate physical model and become costly or unstable in high dimensions [8]. Data-driven surrogates learn unknown or partially known dynamics directly from observations and embed them within Bayesian inversion frameworks. However, the intrinsic stochasticity of complex systems exposes limitations: neural operators such as FNO [33] and Transolver [64] produce only deterministic forecasts and thus cannot quantify uncertainty, while diffusion models [39, 45, 49] struggle to learn physically faithful distributions due to the long Markov chain required to map Gaussian noise into realistic data.
Our contributions.
We address the above limitations by building on recent flow-based stochastic interpolant methods [1, 2, 16], which learn a short-step conditional transition $p(\bm{x}_{k+1}\!\mid\!\bm{x}_{k})$ instead of a global noise-to-data map. Our main contributions are:
  • FlowDAS: We introduce a generative data-assimilation framework that treats a learned stochastic interpolant as the forward model and assimilates new observations on the fly at inference time—without retraining or auxiliary filtering steps.
  • Efficient, interpretable learning: By modeling the transition between adjacent states, FlowDAS trains faster and more stably than long-horizon diffusion bridges, rolls out autoregressively, and provides clearer physical insight into one-step dynamics.
  • Extensive validation: Across the Lorenz-63 system, incompressible Navier–Stokes super-resolution and sparse-sensor tasks, and large-scale SEVIR weather forecasting—where dynamics range from known to partially or fully unknown—FlowDAS consistently outperforms strong baselines in both accuracy and physical plausibility.
2 Preliminaries
2.1 Related Work
Data assimilation alternates between a forecasting step—propagating the current state forward under a dynamical model—and an inverse step that corrects this forecast with new observations.
Model-driven methods
Traditional model-driven approaches include the Kalman filter and its many variants [3]. The Kalman filter provides an optimal minimum-variance estimate when the dynamics are linear and all errors are Gaussian, but its performance degrades once these assumptions are not met. Also, there are some variants for nonlinear settings like the Extended Kalman Filter [47, 43], applying the nonlinear transition to predict states and its Jacobian to the covariance, struggling in discontinuity situations and causing computational expense to calculate the Jacobian. The Unscented Kalman Filters [29] bypass the computation of Jacobian by propagating a set of sample points with the nonlinear transformations and deriving the mean and covariance, making it sensitive to sample point parameters and non-Gaussian problems. For fully nonlinear, non-Gaussian problems, the bootstrap particle filter (BPF) [26] approximates the posterior with a set of weighted particles that are advanced by the dynamics, reweighted by the observation likelihood, and resampled to prevent weight collapse. Although flexible, the BPF requires an exponential number of particles as the state dimension grows, making it impractical for large physical systems [8]. Besides, both the Kalman family and particle filters rely on having accurate knowledge of the underlying dynamical models, a requirement that is often unmet in real-world applications.
Neural operator-based methods
Neural-operator solvers treat forecasting as a deterministic map from the current state to a single future state. The Fourier Neural Operator (FNO; [33]) learns this map in Fourier space, giving fast and accurate predictions, while Transolver ([64]) replaces Fourier layers with a physics-attention Transformer that can effectively capture complex physical correlations. Both models are deterministic and they are not designed for dynamics that are inherently stochastic. Recent work has sought to introduce stochasticity by fitting probabilistic surrogates—such as stochastic processes, graphical models, or stochastic Koopman operators [62, 70]—but these efforts typically focus on low-dimensional settings or assume simple Gaussian statistics, limiting their applicability to high-dimensional, strongly nonlinear systems. Researchers have also adapted neural operator-based models for DA tasks [15, 66, 65]. However, these adaptations are typically tailored to narrow use cases—such as arbitrary-resolution data assimilation—and have not demonstrated robust generalization across diverse real-world scenarios.
Diffusion model-based methods
@@ 2.1 Related Work
Recent works have leveraged score-based methods, in particular diffusion models, for dynamical system modeling. For instance, the Conditional Diffusion Schrödinger Bridge (CDSB) framework extends diffusion models to conditional simulation and demonstrates strong performance on filtering in state-space models [48]. Score-based data assimilation (SDA) is a recent data-driven approach that employs score-based diffusion models to estimate state trajectories in dynamical systems [45]. It bypasses explicit physical modeling by learning the joint distribution of short state trajectory segments (e.g., $12k+$ 1 time steps) via a score network $\bm{s}_{\theta}(\bm{x}_{i-k:i+k})$. By integrating the score network with the observation model in a diffusion posterior sampling (DPS) framework [17], SDA can generate entire state trajectories in zero-shot and non-autoregressive manners. However, SDA captures state transition implicitly from state concatenation, the learned dynamics of SDA may lack physical interpretability, leading to potential inaccuracies. And its posterior approximations may be less reliable in systems where the physical model is highly sensitive, as illustrated by the double well potential problem in Section D.1. Additionally, diffusion models’ non-autoregressive nature makes it hard to deal with long sequence rolling-out problems in high-dimension complex dynamics, as illustrated by the Navier–Stokes scenarios in Figure 3.
@@ 2.2 Stochastic Interpolants
Stochastic interpolants [1, 2, 16] is a generative modeling framework that unifies flow-based and diffusion-based models, which provides a smooth, controlled transition between arbitrary probability densities over a finite time horizon.
Consider a stochastic process $\bm{X}_{s}$ defined over the interval $s\in[0,1]$, which evolves from an initial state $\bm{X}_{0}\sim\pi(\bm{X}_{0})$ to the target state $\bm{X}_{1}\sim q(\bm{X}_{1})$. A stochastic interpolant can be described as:
@@
where $(\bm{X}_{0},\bm{X}_{1})\sim p(\bm{X}_{0},\bm{X}_{1})$. $\bm{W}_{s}$ is a Wiener process for $s\in[0,1]$ introduced after $\bm{X}_{0}$ and $\bm{X}_{1}$ are sampled, ensuring that $\bm{W}_{s}$ is independent of $\bm{X}_{0}$ and $\bm{X}_{1}$. The time-varying coefficients $\alpha_{s},\beta_{s},\sigma_{s}\in C^{1}([0,1])$ satisfy boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=\sigma_{1}=0$, ensuring that $\tilde{\bm{I}}_{0}=\bm{X}_{0}$ and $\tilde{\bm{I}}_{1}=\bm{X}_{1}$, thereby creating a smooth interpolation from $\bm{X}_{0}$ to $\bm{X}_{1}$. Moreover, for all $(s,\bm{X}_{0})\in[0,1]\times\mathbb{R}^{D}$, $\tilde{\bm{I}}_{s}\mid\bm{X}_{0}$ has the same distribution as $\bm{X}_{s}$, which is the solution to the following SDE [16]:
@@
The “velocity” of the interpolant path, $\bm{R}_{s}$, is given by $\bm{R}_{s}=\dot{\alpha}_{s}\bm{X}_{0}+\dot{\beta}_{s}\bm{X}_{1}+\dot{\sigma}_{s}\bm{W}_{s}$. Furthermore, the drift term $\bm{b}_{s}(\bm{X}_{s},\bm{X}_{0})$ is related to the score function $\nabla\log p(\bm{X}_{s}\mid\bm{X}_{0})$:
$\bm{b}_{s}(\bm{X}_{s},\bm{X}_{0})$ $=\frac{\bm{c}_{s}(\bm{X}_{s},\bm{X}_{0})}{\beta_{s}}+\frac{\nabla\log p(\bm{X}_{s}\mid\bm{X}_{0})}{\lambda_{s}\beta_{s}},$ (6)
where $\lambda_{s}$ and $\bm{c}_{s}(\bm{X}_{s},\bm{X}_{0})$ are defined to control the score-based dynamics (Section A.2).
Building on this framework, stochastic interpolants can estimate the transition $p(\bm{x}_{k+1}\mid\bm{x}_{k})$ by evolving a latent path $\bm{X}_{s}$ smoothly from $s=0$ (current state) to $s=1$ (next state) in dynamical systems. It provides a probabilistic yet compact surrogate that is more aligned with the true step-to-step evolution of complex dynamical systems. While stochastic interpolants capture complex system dynamics, they does not yet enforce consistency with measurements.
3 Method